# Multi-Objective min-max Online Convex Optimization

### v1

@@ 6.3 Proof of Theorem 8
Discussion: Compared to the limited information setting considered in Section 2, where gradient information is available, with bandit feedback the regret guarantees suffer for the expected min-max OCO problem, however, the degraded guarantees match the best known guarantees for OGD in the bandit setting. The proof of the result is simple because of the modularity of our algorithm and the regret decomposition that decomposes the total regret into regret resulting from Hedge and OGD, and the fact that only regret for OGD changes with bandit feedback.
7 Experimental Results
In this section, we present numerical results on the expected min-max static regret of Algorithm 1 under relevant settings, that are well motivated from both theoretical as well applications point of view. In all the figures, we plot the expected min-max static regret as a function of time horizon $T$.
7.1 Basic Results
We begin our experimental results by considering the input model when all functions $f_{t}^{k}$ for $k=1,\dots,K,t=1,\dots,T$ are random linear functions in $d=10$ dimensions, where each coefficient is distributed uniformly between $[0,1]$. Random linear functions are important since they were sufficient to find lower bounds on the regret of any online algorithm for OCO Hazan (2019). In Fig. 1, we plot the expected min-max static regret of Algorithm 1 as a function of $K$ and $T$ for random linear functions, where as expected, we observe that the regret scales as $O(\sqrt{T\log K})$.
Next, we consider $f_{t}^{k}$’s as strongly convex functions, where each $f_{t}^{k}$ is independently and identically selected from a set of $\{(x-a)^{2},a=[-9,-8,\dots,0,1,\dots,9]\}$. Recall that for OCO with strongly convex functions, the step size $\eta$ needed for OGD to get a regret guarantee of $O(\log T)$ scales as $\eta=1/t$ and not $\eta=1/\sqrt{t}$ Hazan (2019). Thus, for this simulation of Algorithm 1333In contrast to $\eta_{x,t}=O(1/\sqrt{t})$ as needed to prove Theorem 4. we use $\eta_{\theta,t}=O(1/\sqrt{t})$ and $\eta_{x,t}=O(1/t)$.  [^4] We have checked that other combinations only increase the regret. As pointed out earlier, we are unable to improve the expected min-max static regret guarantee for Algorithm 1 when all functions $f_{t}^{k}$’s are strongly convex. We see that in fact that is true for Algorithm 1 even in simulations as shown in Fig. 2, where for $K=1$, which is the same as OCO (with $\eta_{x,t}=O(1/t)$) the regret scales as $\log(T)$, while for larger values of $K$, the expected min-max static regret scales as $O(\sqrt{T})$ when $\eta_{\theta,t}=O(1/\sqrt{t})$ and $\eta_{x,t}=O(1/t)$. Thus, if indeed expected min-max static regret of $O(\log T)$ is achievable then Algorithm 1 needs to be conceptually modified.
050501001001501502002002502503003000202040406060Time ($t$)Average Cumulative RegretExpected min-max static regret for random linear functionsK=1K=1K=10K=10K=50K=50K=100K=100
Figure 1: Expected min-max static regret of Algorithm 1 when $f_{t}^{k}$ are random linear functions for different values of $K$.
0505010010015015020020025025030030002002004004006006008008001,0001{,} [^0] Time Horizon ($T$)Average RegretExpected min-max static regret for quadratic loss functionsK=1K=1K=10K=10K=50K=50K=100K=100
Figure 2: Expected min-max static regret of Algorithm 1 when $f_{t}^{k}$ are random quadratic functions for different values of $K$.
7.2 GlobalExpertsProblem
In this section, we compare the performance of Algorithm 1 and MULTI Even-Dar et al. (2009) for the GlobalExpertsProblem with the i.i.d. input. For that purpose, we consider that the loss of each expert at time $t$ is drawn independently from a uniform distribution:
$\ell_{t,k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathsf{Uniform}(a,b),\quad t=1,2,\dots,T,\;k=1,2,\dots,K,$
where $a,b\in[0,1]$ are fixed parameters defining the range of losses. In our experiments, we set $a=0.2$ and $b=0.8$ to ensure non-trivial loss variation while avoiding extreme loss values.
For this model, we plot the expected min-max static regret performance of Algorithm 1 and MULTI in Fig. 3, and see that both perform similarly, even though the regret guarantee we obtained for Algorithm 1 in i.i.d. case is slightly better than MULTI (which is valid even with the adversarial input). It appears that MULTI has better performance than its guarantee Even-Dar et al. (2009). However, MULTI is computationally far more intensive than Algorithm 1 since it uses the $K=2$ primitive recursively, and moreover, it needs to know the time horizon $T$ in contrast to Algorithm 1.
1,0001{,}0002,0002{,}0004,0004{,}0008,0008{,}0000551010151520202525303035354040K=2K=2K=4K=4K=16K=16K=32K=32K=64K=64Time horizon TTAverage RegretMULTIAlgorithm 1
Figure 3: Expected min-max static regret of MULTI Even-Dar et al. (2009) and Algorithm 1 for different numbers of experts $K$ and time horizons $T$ for the fair classification example.
7.3 Application of min-max formulation to Fair Classification Setting
Finally, in this subsection, we simulate the i.i.d. min-max OCO problem for a fair classification setting.
Data generation.
We consider $K$ groups, each corresponding to a convex loss function sequence. For each group $k\in[K]$, we sample a hidden parameter $w_{k}^{\star}\sim\mathcal{N}(0,I_{d})$. At each round $t=1,\dots,T$, for each group $k$ we draw $m$ feature vectors
$z_{t,k,i}\sim\mathcal{N}(\mu_{k},\sigma^{2}I_{d}),\quad i=1,\dots,m,$
with $\mu_{k}$ drawn once per group. Labels are generated as
$y_{t,k,i}\sim\mathrm{Bernoulli}\!\left(\sigma(\langle w_{k}^{\star},z_{t,k,i}\rangle)\right),$
where $\sigma(u)=1/(1+e^{-u})$ is the sigmoid function.
Loss functions.
The per-group convex loss at round $t$ is the averaged logistic loss with $L_{2}$ regularization:
$f_{t}^{k}(x)\;=\;\frac{1}{m}\sum_{i=1}^{m}\log\!\Big(1+\exp(-y_{t,k,i}\,\langle x,z_{t,k,i}\rangle)\Big)\;+\;\kappa\|x\|_{2}^{2}.$
For this application, we compare the performance of following algorithms.
  • Algorithm 1 (OGD+Hedge)
  • OGD on average loss: projected OGD on $\tfrac{1}{K}\sum_{k=1}^{K}f_{t}^{k}(x)$.
  • Follow-the-Regularized-Leader (FTRL): $x_{t}=\arg\min_{x\in\mathcal{X}}\left\{\max_{k}\sum_{s=1}^{t-1}f_{s}^{k}(x)+\mathcal{R}(x)\right\}$.
Parameters.
We use $d=20$, $K=10$, mini-batch size $m=50$, diameter $D=5$, and regularization $\kappa=10^{-3}$. Step sizes are set as $\eta_{x,t}=D/(G\sqrt{t})$ for OGD (with $G$ an upper bound on gradient norms) and $\eta_{\theta,t}=\sqrt{\ln K}/(B\sqrt{t})$ for Hedge, with $B$ an upper bound on per-round losses.
In Fig. 4, we plot the expected min-max static regret ${\bar{R}}_{T}$ for all the three algorithms. We see that all algorithms have sub-linear regret, and Algorithm 1 has larger regret than FTRL but better than average OGD. However, recall that neither FTRL and average OGD have any theoretical guarantees on their expected min-max static regret. Next, in Fig. 5, we will consider an input where the index of the most challenging sequence changes over time and show that Algorithm 1 outperforms FTRL in that case.
10010020020030030040040002020404060608080Time Horizon ($T$)Average Regret vs. Offline OptimalExpected min-max static regret for $K=10$ for the fair classification problemAlgorithm 1 OGD+HedgeAvg-OGDFTL
Figure 4: Expected min-max static regret comparison for $K=10$.
Non-stationary input
Next, we change the input described in Section 7.3 to model a highly non‑stationary environment specifically designed to challenge static optimization strategies such as FTRL and highlight the adaptability of Algorithm 1 that uses OGD+Hedge. The detailed changes with respect to Section 7.3 are as follows:
  • We consider $K=3$ separate objectives, each with its own logistic loss function defined by fixed but distinct parameter vectors $w_{\text{star}}[k]$ and mean feature shifts $\mu[k]$.
  • Every switch_interval steps (here set to 100), one objective becomes significantly harder by adding a large mean shift ($\texttt{shift\_magnitude}=5.0$) to its feature distribution. This creates abrupt changes in which objective dominates the cumulative loss objective function.
  • Data generation: For each round $t$, and for each objective $k$, a feature matrix $Z$ of size $m\times d$ is generated from a Gaussian distribution centered at the shifted mean. Labels $y$ are generated from a logistic model using $w_{\text{star}}[k]$ (is a hidden “true” parameter vector for objective $k$. It defines how that objective behaves.), with additional noise introduced via random sampling.
We clearly see in Fig. 5 that Algorithm 1 (OGD+Hedge) adapts better to changing conditions than FTRL, since worst‑case objective changes over time, making past loss information misleading for FTRL.
1010505010010030030050050002020404060608080Time Horizon ($T$)Average Regret (Avg. over 10 Seeds)Expected min-max static regret in switching environment for K=3K=3Algorithm 1 OGD+HedgeFTRL
Figure 5: Expected min-max static regret for Algorithm 1 vs. FTRL in a piecewise-stationary environment.
@@ 8 Conclusions
In this paper, we have made a basic advance in the area of multi-objective OCO by formulating the min-max static regret minimization problem, that requires an algorithm to closely track multiple loss function sequences simultaneously and at all times. Because of analytical challenges with the adversarial input, we considered the i.i.d. input setting, for which we proposed a simple algorithm that used a combination of the well known Hedge and OGD algorithm. Using an insightful regret decomposition, which made the analysis modular, we derived a $O(\sqrt{T\log K})$ regret bound on the proposed algorithm, which is shown to be best possible via simulations. We also extended the results to the bandit information setting. Certain questions that remain open are: i) what is the best regret possible with the adversarial input, ii) how to handle the case when all loss functions are strongly convex, and iii) handling constraints like COCO.
Appendix A Proof of Lemma 5

### v3

@@ 7 Extension to Bandit Feedback
Discussion: Compared to the limited information setting considered in Section 3, where gradient information is available, with bandit feedback the expected min-max regret guarantees suffer, however, the degraded guarantees match the best known guarantees for OGD in the bandit setting. The proof of the result is simple because of the modularity of our algorithm and the regret decomposition that decomposes the total regret into regret resulting from Hedge and OGD, and the fact that only regret for OGD changes with bandit feedback.
8 Binary Prediction Problem
In this section, we consider the min-max universal prediction of binary sequences, where at each time $t$, two binary sequences $b_{t}^{1},b_{t}^{2}\in\{0,1\}$ are revealed after an online algorithm $\mathcal{A}$ (predicts) chooses an action $x_{t}\in\{0,1\}$, and cost of $\mathcal{A}$ is
$C^{b}_{\mathcal{A}}\;=\;\max\left\{\sum_{t=1}^{T}\mathbf{1}\{x_{t}\neq b_{t}^{1}\},\;\sum_{t=1}^{T}\mathbf{1}\{x_{t}\neq b_{t}^{2}\}\right\}.$
Similar to (2),
$C^{b}_{\mathsf{OPT}}\;=\;\min_{x\in\{0,1\}}\max\Big\{\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{1}\},\;\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{2}\}\Big\},$
and the regret is defined as $R^{b}_{\mathcal{A}}(T)=\max_{b_{t}^{1},b_{t}^{2}}\{C^{b}_{\mathcal{A}}-C^{b}_{\mathsf{OPT}}\}.$
Similar to Theorem 4, we show that the min-max static regret $R^{b}_{\mathcal{A}}(T)$ for any online algorithm is $\Omega(T)$ with adversarial input in contrast to the classical setting Cover (1966); Feder et al. (2002) where with a single binary sequence the regret is $O(\sqrt{T})$ with adversarial input.
Theorem 10.
The min-max static regret $R^{b}_{\mathcal{A}}(T)$ for any online algorithm is $\Omega(T)$.
The proof of Theorem 10 can be found in Appendix K. Thus, once again we consider the stochastic i.i.d. input as follows.
8.1 $(b_{t}^{1},b_{t}^{2})$ i.i.d. across time
In this section, we consider that $(b_{t}^{1},b_{t}^{2})$ are generated from an unknown distribution with i.i.d. across time. Under this setting, we can directly use Theorem 5 to show that Algorithm 1 can be used to get a true min-max static regret bound of ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(\sqrt{T\log T})$, see details in Appendix L.
Next, we show that in this binary case, we can do better than Algorithm 1 and get true min-max static regret bound ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(1)$ without involving the averaged-out benchmark ${\bar{C}}_{\mathsf{OPT}}^{b}$ that is defined similar to (7).
The main challenge in this binary min-max prediction setting arises in disagreement regimes, where the two sequences $\{b_{t}^{1}\}_{t=1}^{T},\{b_{t}^{2}\}_{t=1}^{T}$ favor opposite actions. In such cases, simply following a majority or resolving ties arbitrarily can lead to persistent oscillations, since each sequence individually advocates a different decision. To achieve low expected min-max static regret, an algorithm must determine which of these conflicting preferences induces the larger worst-case cost and commit to that action.
To address this we propose Adaptive Stronger-Bias Double Majority (ASB-DM) algorithm (Algorithm 4 in Appendix M) that exploits the strength of the statistical evidence, rather than only its direction. When the empirical preferences of the two sequences disagree, ASB-DM aggregates the evidence by comparing the combined bias toward the two actions and selects the one with the smaller inferred worst-case loss. As the empirical estimates concentrate, this stronger-bias rule allows the algorithm to reliably distinguish the optimal action from the suboptimal one, ensuring that the decision stabilizes and the total number of mistakes remains finite.
Theorem 11.
Assume the pairs $(b_{t}^{1},b_{t}^{2})$ are generated i.i.d. across time $t$ from a joint distribution with marginals $b_{t}^{1}\sim\mathrm{Ber}(q_{1})$ and $b_{t}^{2}\sim\mathrm{Ber}(q_{2})$. If the preferences are asymmetric (i.e., $q_{1},q_{2}\neq 1/2$ and $q_{1}+q_{2}\neq 1$), then the true min-max regret of Algorithm 4 satisfies ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(1)$.
Proof of Theorem 11 is provided in Appendix M. Note that extension to the case when there are $K\geq 2$ sequences $b_{t}^{k},k=1,\dots,K$ is immediate with $O(1)$ term scaling linearly with $K$.
Remark 8 (Singularity at $q_{1}+q_{2}=1$).
The necessity of the condition $q_{1}+q_{2}\neq 1$ is best illustrated by the example where $q_{1}=0.4$ and $q_{2}=0.6$. To determine the static optimal action, we compare the worst-case expected costs: action 0 incurs $0.6T$ (driven by Sequence 2), which is identical to the $0.6T$ cost incurred by action $1$ (driven by Sequence 1). Since the worst-case costs are identical, the static optimal action is not unique. At this singularity, any online algorithm faces an inherent lack of statistical separation between the actions. Because the expected costs are perfectly balanced, empirical observations will fluctuate around the true parameters due to random sampling noise. Without a strictly positive gap to stabilize the decision process, any algorithm will be forced to switch between actions repeatedly—effectively behaving as a symmetric random walk—leading to an accumulation of regret at a rate of $\Omega(\sqrt{T})$.
@@ 9 Conclusions
In this paper, we have made a basic advance in the area of multi-objective OCO by formulating the min-max static regret minimization problem, that requires an algorithm to closely track multiple loss function sequences simultaneously and at all times. Because of the lower bound on regret that scales linearly with the time horizon under an adversarial input, we considered the i.i.d. input setting, for which we proposed a simple algorithm that used a combination of the well-known Hedge and OGD algorithm. Using an insightful regret decomposition, which made the analysis modular, we derived an $O(\sqrt{T\log(TK)})$ regret bound on the proposed algorithm. We also extended the results to the bandit information setting. Certain questions that remain open are: i) how to handle the case when all loss functions are strongly convex, and ii) handling constraints like COCO.
We also generalized the classical problem of universal prediction of binary sequences to the case when there are multiple sequences. Even for this binary case, we showed that the min-max static regret of any online algorithm scales linearly with the time horizon, exposing the inherent difficulty of the min-max setting. With stochastic input where $(b_{t}^{1},b_{t}^{2})$ are i.i.d. across time, we proposed an algorithm that chooses its action on the strength of statistical evidence and showed that it has $O(1)$ expected min-max static regret.
Appendix A Applications of Multi-objective OCO and Open Questions
A.1 Applications
The min-max approach is also well suited for modern networked systems—such as data centers, wireless edge networks, and shared cloud infrastructures—where online decisions must balance performance across multiple service classes, users, or flows whose objectives (throughput, latency, energy) often conflict. The min-max regret criterion provides a principled way to ensure worst-case fairness by minimizing the largest cumulative loss among competing entities, which is crucial in fair resource allocation [43], fair allocation in clouds [19], fair assignment of reusable resources [28] congestion-controlled transport [24, 3], and dynamic resource provisioning in data centers [45]. Unlike average-regret formulations that can conceal systematic bias, minimizing min-max regret explicitly guards against persistent under-allocation to any group or flow—an increasingly important property in multi-tenant and QoS-sensitive networks.
Multi-objective OCO has also been studied from an applications perspective, e.g. [30] studied the problem of multi-period multi-objective online ride-matching problem, where the multiple objectives included platform revenue, assigning drivers with higher service quality, minimizing the distance to the customer etc. [30] also used the COCO-type reformulation of multi-objective OCO for deriving theoretical guarantees. Other applications of multi-objective OCO include multi-task learning [38], multiple target tracking [33], portfolio selection [12] and safe reinforcement learning [10], where decision-makers must make sequential decisions under uncertainty, balancing objectives such as cost, delivery time, and reliability without full knowledge of future demand or disruptions.
A.2 Open Questions
    and the corresponding regret and CCV are
    respectively. It turns out that our proposed algorithm cannot be married readily with the prior best known approaches for COCO [41, 48, 42] for deriving simultaneous bounds on regret and CCV. Thus, how to solve the min-max OCO with constraints remains an interesting open problem.
Appendix B Related Work on COCO
Because of its widespread applications, multi-objective OCO has been an object of immense interest, however, with limited success. For example, [32] wanted to study the multi-objective OCO and stated the difficulty as “A fundamental difference between single- and multi-objective optimization is that for the latter it is not obvious how to evaluate the optimization quality”. Further, it said “Since it is impossible to simultaneously minimize multiple loss functions and in order to avoid complications caused by handling more than one objective, we choose one function as the objective and try to bound other objectives by appropriate thresholds. This work gave rise to the now popular problem of constrained online convex optimization (COCO), where the problem is transformed into a constrained optimization problem with one objective and one or more constraint functions, that is described next.
B.0.1 COCO
$\textrm{CCV}_{\mathcal{A}}(T)\equiv\sum_{t=1}^{T}\max(g_{t}(x_{t}),0).$
B.0.2 Prior Work on COCO
(A) i.i.d. loss and constraints: COCO with independent and identically distributed (i.i.d.) loss functions $f_{t}$’s and constraint functions $g_{t}$’s was studied in [32] while its extension to stationary and ergodic processes can be found in [46], where both derived an algorithm with regret $O(\sqrt{T})$ and CCV $O(T^{3/4})$.
(B) Time-invariant constraints: COCO with time-invariant constraints, i.e., $g_{t}=g,\forall\ t$ [53, 21, 31, 50] has been considered extensively, where functions $g$ are assumed to be known to the algorithm a priori. The best known results in this case give static regret of $O(T^{\max\{c,1-c\}})$ and CCV of $O(T^{(1-c)/2})$ for any $c\in(0,1)$. The algorithm is allowed to take actions that are infeasible at any time to avoid the costly projection step of the vanilla projected OGD algorithm and the main objective was to design an efficient algorithm with a small regret and CCV while avoiding the explicit projection step.
(C) Time-varying constraints: The more difficult question is solving COCO problem when the constraint functions, i.e., $g_{t}$’s, change arbitrarily with time $t$. One popular algorithm for solving COCO considered a Lagrangian function optimization that is updated using the primal and dual variables [52, 44, 51]. Alternatively, [35] and [26] used the drift-plus-penalty (DPP) framework [34] to solve the COCO, but which needed additional assumption, e.g. the Slater’s condition in [35] and with weaker form of the feasibility assumption in [35].
[18] obtained bounds similar to [35] but without assuming Slater’s condition. Very recently, the state of the art guarantees on simultaneous bounds on regret $O(\sqrt{T})$ and CCV $O(\sqrt{T}\log T)$ for COCO were derived in [41] with a very simple algorithm that combines the loss function at time $t$ and the CCV accrued till time $t$ in a single loss function, and then executes the online gradient descent (OGD) algorithm on the single loss function with an adaptive step-size. Most recently [42] showed lower CCV is possible with increased regret compared to [41] while [48] obtained instance dependent CCV bounds which can be $O(1)$ for specific cases while achieving regret of $O(\sqrt{T})$, respectively. The COCO problem has been considered in the dynamic setting as well [8, 4, 47, 29] where the benchmark $x^{\star}$ in (14) is replaced by $x_{t}^{\star}$ ($x_{t}^{\star}=\arg\min_{x}f_{t}(x)$) that is also allowed to change its actions over time.
Remark 9.
Appendix C Proof of Theorem 4
Proof.
Let the decision set be $\mathcal{X}=[-1,1]$. The adversary constructs two inputs $\sigma_{1}$ and $\sigma_{2}$ with common prefix until $t=T/2$ as follows. For both inputs $\sigma_{1}$ and $\sigma_{2}$, for rounds $t=1,\dots,T/2$,
$f_{t}(x)=-x,\qquad g_{t}(x)=x.$
Let $x_{t}$ be the actions of an online algorithm $\mathcal{A}$, and let
${\bar{x}}:=\sum_{t=1}^{T/2}x_{t}$
denote the cumulative action of $\mathcal{A}$ during the prefix. After observing $\bar{x}$}, the adversary completes the input as follows:
  • If ${\bar{x}}<0$, the adversary selects input $\sigma_{1}$.
  • If ${\bar{x}}\geq 0$, the adversary selects input $\sigma_{2}$, where the respective suffixes for $t=T/2+1,\dots,T$ are defined as follows.
Suffix of $\sigma_{1}$.
For $t=T/2+1,\dots,T$,
$f_{t}(x)=1,\qquad g_{t}(x)=0.$
Suffix of $\sigma_{2}$.
For $t=T/2+1,\dots,T$,
$f_{t}(x)=0,\qquad g_{t}(x)=1.$
For $\sigma_{1}$, the cumulative losses for a fixed $x\in[-1,1]$ are
$\sum_{t=1}^{T}f_{t}(x)=-\frac{T}{2}x+\frac{T}{2},\qquad\sum_{t=1}^{T}g_{t}(x)=\frac{T}{2}x.$
The two quantities are equal at $x=\tfrac{1}{2}$, yielding
$C_{\mathsf{OPT}}(\sigma_{1})=\frac{T}{4}.$
Similarly, for $\sigma_{2}$,
$\sum_{t=1}^{T}f_{t}(x)=-\frac{T}{2}x,\qquad\sum_{t=1}^{T}g_{t}(x)=\frac{T}{2}x+\frac{T}{2},$
which are equal at $x=-\tfrac{1}{2}$, giving
$C_{\mathsf{OPT}}(\sigma_{2})=\frac{T}{4}.$
Thus, in both cases,
$C_{\mathsf{OPT}}(\sigma_{i})=\frac{T}{4},\qquad i=1,2.$
We now lower bound the cost incurred by $\mathcal{A}$.
Case 1: ${\bar{x}}<0$ (input $\sigma_{1}$).
$\sum_{t=1}^{T}f_{t}(x_{t})=-\sum_{t=1}^{T/2}x_{t}+\frac{T}{2}=\frac{T}{2}-{\bar{x}},$
$\sum_{t=1}^{T}g_{t}(x_{t})=\sum_{t=1}^{T/2}x_{t}={\bar{x}}.$
Hence,
$C_{\mathcal{A}}(\sigma_{1})=\max\Bigl\{\frac{T}{2}-{\bar{x}},\;{\bar{x}}\Bigr\}\geq\frac{T}{2},$
since ${\bar{x}}<0$.
Case 2: ${\bar{x}}\geq 0$ (input $\sigma_{2}$).
$\sum_{t=1}^{T}f_{t}(x_{t})=-\sum_{t=1}^{T/2}x_{t}=-{\bar{x}},$
$\sum_{t=1}^{T}g_{t}(x_{t})=\sum_{t=1}^{T/2}x_{t}+\frac{T}{2}.$
Thus,
$C_{\mathcal{A}}(\sigma_{2})=\max\Bigl\{-{\bar{x}},\;\frac{T}{2}+{\bar{x}}\Bigr\}\;\geq\;\frac{T}{2},$
since ${\bar{x}}\geq 0$.
Thus,
$\mathrm{Regret}_{\mathcal{A}}(T)=C_{\mathcal{A}}(\sigma_{i})-C_{\mathsf{OPT}}(\sigma_{i})\geq\frac{T}{4}$
for at least one $i=1,2$. ∎
Appendix D Connecting the two benchmarks ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ and ${\bar{C}}_{\mathsf{OPT}}$ (7)
Without loss of generality, let $\chi\subseteq[0,1]^{d}$. We will invoke Assumptions 1 and 2 throughout without repeatedly mentioning it. Thus, for each $t=1,\dots,T$ and $x\in[0,1]^{d}$, we have
$\lambda_{t}(x)=(f_{t}(x),g_{t}(x))\in[0,B]^{2},$
where $\{\lambda_{t}(\cdot)\}_{t=1}^{T}$ are i.i.d. across $t$. Since $f_{t}(x),g_{t}(x)$ is $G$-Lipschitz, each coordinate of $\lambda_{t}$ is GG–Lipschitz in $x$. Let the empirical mean be
$\hat{\mu}_{T}(x)=\frac{1}{T}\sum_{t=1}^{T}\lambda_{t}(x),$
while as before
$\mu(x)=\mathbb{E}\{\lambda_{t}(x)\}.$
Let $\Phi=\min_{x}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x).$ Then from (2) and (7)
$C_{\mathsf{OPT}}=\min_{x}\max\Big\{\sum_{t}f_{t}(x),\,\sum_{t}g_{t}(x)\Big\},\qquad{\bar{C}}_{\mathsf{OPT}}=T\Phi.$
Let $\phi(v):=\min_{x}\max_{\theta\in\Delta_{2}}\theta^{\top}v(x)=\min_{x}\max\{v_{1}(x),v_{2}(x)\}$ for $v:[0,1]^{d}\to\mathbb{R}^{2}$.
Lemma 12 (1–Lipschitz property of $\phi$).
For any $v,w:[0,1]^{d}\to\mathbb{R}^{2}$,
$\big|\phi(v)-\phi(w)\big|\;\leq\;\sup_{x\in[0,1]^{d}}\ \|v(x)-w(x)\|_{\infty}.$
Proof.
For each $x\in[0,1]^{d}$ define
$h_{v}(x)\;:=\;\max_{\theta\in\Delta_{2}}\theta^{\top}v(x)\;=\;\|v(x)\|_{\infty}.$
Then $\phi(v)=\min_{x}h_{v}(x)$.
First, note the elementary inequality valid for any $a,b\in\mathbb{R}^{2}$:
$\max_{k}a_{k}-\max_{k}b_{k}\;\leq\;\max_{k}(a_{k}-b_{k})\;=\;\|a-b\|_{\infty}.$
Using ($\ast$) pointwise with $a=v(x)$ and $b=w(x)$ gives
$h_{v}(x)-h_{w}(x)\;=\;\sup_{k}v_{k}(x)-\sup_{k}w_{k}(x)\;\leq\;\sup_{k}\big(v_{k}(x)-w_{k}(x)\big)\;=\;\|v(x)-w(x)\|_{\infty}.$
Thus, we get
$\phi(v)-\phi(w)\;=\;\inf_{x}h_{v}(x)-\inf_{x}h_{w}(x)\;\leq\;\sup_{x}\big(h_{v}(x)-h_{w}(x)\big)\;\leq\;\sup_{x}\|v(x)-w(x)\|_{\infty}.$
Swapping the roles of $v$ and $w$ gives the reverse inequality $\phi(w)-\phi(v)\leq\sup_{x}\|v(x)-w(x)\|_{\infty}$. Combining the two displays proves
$\big|\phi(v)-\phi(w)\big|\;\leq\;\sup_{x\in[0,1]^{d}}\ \|v(x)-w(x)\|_{\infty}.$
Hence, from Lemma 12
$\Big|\frac{C_{\mathsf{OPT}}}{T}-\Phi\Big|=|\phi(\hat{\mu}_{T})-\phi(\mu)|\leq\sup_{x\in[0,1]^{d}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}.$ (14)
Let $\mathcal{N}_{r}$ be an $\ell_{\infty}$ rr–net of $0,1]^{d} [$49], then $|\mathcal{N}_{r}|\leq(3/r)^{d}$ Lemma 5.2 [49]. Using the GG–Lipschitzness of $f_{t},g_{t}$, from Lemma 5.2 [49]
$\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\leq\max_{x\in\mathcal{N}_{r}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}+2G\,r.$ (15)
For each fixed $x\in\mathcal{N}_{r}$ and coordinate $j\in\{1,2\}$, Hoeffding’s inequality for [0,B][0,B]–bounded variables gives
$\Pr\!\left(\big|\hat{\mu}_{T}^{(j)}(x)-\mu^{(j)}(x)\big|>\gamma\right)\leq 2\exp\!\left(-\frac{2T\gamma^{2}}{B^{2}}\right).$
A union bound over $\mathcal{N}_{r}$ and the two coordinates yields
$\Pr\!\left(\max_{x\in\mathcal{N}_{r}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}>\gamma\right)\leq 4|\mathcal{N}_{r}|\exp\!\left(-\frac{2T\gamma^{2}}{B^{2}}\right).$ (16)
Integrating the tail probability in (16) gives
$\mathbb{E}\!\left\{\max_{x\in\mathcal{N}_{r}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}$ $\leq B\sqrt{\frac{\log(4|\mathcal{N}_{r}|)}{2T}}+\frac{\sqrt{\pi}\,B}{2\sqrt{2T}}.$ (17)
Combining (15) and (17) yields
$\mathbb{E}\!\left\{\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}\leq B\sqrt{\frac{\log(4|\mathcal{N}_{r}|)}{2T}}+\frac{\sqrt{\pi}\,B}{2\sqrt{2T}}+2G\,r.$ (18)
Take $r=T^{-1/2}$, so $|\mathcal{N}_{r}|\leq(3\sqrt{T})^{d}$ and $\log(4|\mathcal{N}_{r}|)\leq c_{0}d+\tfrac{d}{2}\log T$ with $c_{0}=\log 12$. Then (18) gives
$\mathbb{E}\!\left\{\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}\leq C_{1}\,B\sqrt{\frac{d\log T}{T}}+C_{2}\frac{1}{\sqrt{T}}+C_{3}\,\frac{G}{\sqrt{T}},$
for constants $C_{1},C_{2},C_{3}$.
Taking expectations in (14) and multiplying by $T$,
$\mathbb{E}\left\{\big|C_{\mathsf{OPT}}\}-T\Phi\big|\right\}\leq T\,\mathbb{E}\!\left\{\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}\leq C_{1}\,B\,\sqrt{d\,T\,\log T}+C_{2}\sqrt{T}+C_{3}\,G\,\sqrt{T}.$ (19)
Remark 10.
If there are $K$ functions $f_{t}^{k}$, $k=1,\dots,K$, instead of just two $f_{t},g_{t}$, the bound (19) changes by an additive term $O(\sqrt{T\log K})$, resulting in the final bound of $O(\sqrt{T\log(TK)})$.
Appendix E Proof of Lemma 6